40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  ATF eLog, Page 1 of 54  Not logged in ELOG logo
New entries since:Wed Dec 31 16:00:00 1969
ID Date Author Type Category Subject
  2697   Fri Nov 19 14:01:40 2021 Stephen, RadhikaDailyProgressCryo vacuum chamberRadiative Cooling of Si Mass, with better shield emissivity

Posting cooldown trends for 11/16-11/19 [Attachment 1]. For reference, Attachment 2 shows cooldown data from the last time the Si mass was in the chamber.

Attachment 1: cooldown_1116-1119.pdf
cooldown_1116-1119.pdf
Attachment 2: OLD_cooldown_728.pdf
OLD_cooldown_728.pdf
  2696   Fri Nov 19 09:36:52 2021 AnchalMiscEquipment transferBorrowing two 2" optics

I came and looked around for a Y1S HR coated 2" optic. I found two ATFilms labeled optics with information of only their substrate. The coating info is encoded in the run# but I could not find a place for what it means. So I'm gonna take these and measure the transmission in 40m.

  • 2"x.375" thk FS, HR/AR @ 1064, Run #'s: V6-704 & V6-705
  • 2" x 0.375" PL/PL Run #'s: V2-2239 & V2-2242
Attachment 1: PXL_20211119_174146457.jpg
PXL_20211119_174146457.jpg
  2695   Fri Nov 12 14:31:38 2021 Stephen, RadhikaDailyProgressCryo vacuum chamberRadiative Cooling of Si Mass, with better shield emissivity

In this phase, we are working toward improving our setup with a rigid copper bar, and obtaining a new data point for our radiative cooling thermal models for a suspended silicon mass. Since the past cooling runs of a silicon test mass did not yet incorporate aquadag-painted shields, we wanted to obtain a new data point in the model (in other words, we painted the shields in QIL/2645, but the next test was a PD measurement, so this is the first silicon test mass measurment after shields were painted). The improvement to the thermal linkage, now using a rigid copper bar with higher conductivity (ref. QIL/2666), is a second variable being changed simultaneously in the spirit of improving the cooldown time.

Refer to the prior post (QIL/2694) for the bulk of the blow-by-blow of configuring the chamber to use the rigid copper bar linkage. This post will describe the mounting of the Si mass, and the pump down and cool down.

  • The silicon mass with Aquadag barrel was dropped into the existing frame, with the previous wire arrangement and with no particular requirement on position or orientation (just best effort centering and leveling). Adjustments were done chamberside as access was easier.
  • The frame was lifted into the chamber, with the hanging mass supported by auxiliary fingers, and placed in an available area. Since conductive cooling was not a dominant mode of heat transfer in this setup (ref. QIL/2647), clamping to the baseplate was simply a single dog clamp on each foot of the frame.
  • The cigarette paper was cryovarnished to the surface in the bare central position. Once the cryovarnish was set, the RTD was cryovarnished to the cigarette paper pad. No  strain relief or thermal anchoring considerations were implemented. RTD continuity was verified.
  • Lids were bolted down and shields were finalized (avoiding shorting to copper bar, making sure foil drapes covering apertures were well positioned, etc.
  • Vacuum pumps on at ~3 pm, cryocooler on at 3:30 pm. At 4 pm, things are still looking good!

Closeout photos will be posted to the QIL Cryo Vacuum Chamber photo album.

Attachment 1: IMG_2738.jpeg
IMG_2738.jpeg
Attachment 2: IMG_2741.jpeg
IMG_2741.jpeg
Attachment 3: IMG_2742.jpeg
IMG_2742.jpeg
Attachment 4: IMG_2745.jpeg
IMG_2745.jpeg
Attachment 5: IMG_2747.jpeg
IMG_2747.jpeg
  2694   Fri Nov 12 14:21:32 2021 Stephen, RadhikaDailyProgressCryo vacuum chamberUpgrade to Rigid Copper Bar, and assorted transitions from PD testing

WIP

This post describes upgrade efforts from 10 - 16 November, with the following goals:

 - introducing a solid copper bar thermal linkage

 - shifting the setup away from PD testing

 - preparing for the next test (radiative cooling of Si)

Here are some highlights of the effort:

  • While removing the shields we found contaminants had plated near the line of sight surfaces at the optical window and the electrical feedthrough [Attachment 1]. The film was removed by IPA wipe [Attachment 2] and was not evident in any other location (presumably these were the cold surfaces in line of sight, so they received the most contaminant, but there may be a thinner deposition throughout!).
  • In order to install the copper rod, we needed to cut out slots in the outer shield and inner shield. We used a reciprocating saw and held the piece stiffly on a table [Attachment 3] [Attachment 4].
    • We tried to use large snips, but that failed to provide enough cutting force, especially where it was necessary to use the tip to access into the flanged corner. We also damaged a few of the electrical leads to the feedthrough (formerly used to wire the OSEMs) - this will require attention at next opportunity.
  • The copper rod itself was found to have a few issues:
    • Outer surfaces were somewhat tarnished and greasy, so an 80 grit aluminum oxide paper was used to clean up all surfaces [Attachment 5]. The surfaces were then wiped with IPA using Alpha wipes.
    • Slot width was matching the long thermal strap instead of the short thermal strap, so a drill press was used to add cut away the correct areas of one pair of holes [Attachment 6]. Later, this modification required larger washers in the stack under the bolt head.
    • The cold head had a larger OD than designed, so we were unable to use the corner holes to mount the RTD. There was not enough space at the corners to host the nut or washer. Instead one of the bolt pattern holes was used to host the RTD stack [Attachment 7]
  • The installation of the copper rod required the following steps:
    • We documented the previous state of the table [Attachment 8].
    • We removed the inner shield, the outer shield, and the old thermal linkage.
      • To access the cold head side of the linkage, we had to remove the lower conflat of the T.
    • We installed a 6" conflat flange, double faced with 4" bore (Lesker p/n DFF600X400), between the cryocooler and the T. This spacer raised the bottom face of the cold head to the correct height, so that the rigid copper bar runs approximately through the center of the shield apertures.
      • This required an order of new bolts with 3" length, to squeeze the three flanges (cryocooler, spacer, T) that are now stacked together.
    • We bolted the rigid copper bar to the coldhead, yawing the cryocooler to match the conflat bolt hole orientation while also pointing the copper bar down the axis of the arm.
      • This used new vented screws (UC Components p/n C-412-A) which are also silver plated, except in the location borrowed for the RTD stack as mentioned above [Attachment 7]. The RTD stack included a nut to adjust spring compression independently from screw threading.
      • Apiezon N thermal grease was applied on both surfaces to improve thermal conductivity across this joint.
      • We initially forgot to reinstall the mylar sheet radiation shielding that had been removed from the area around the cold head and around the linkage. This required that we reopen the bottom conflat to install the coldhead mitten, and that we pitch the aluminum shields away from the rigid bar to allow the mylar sheet to be inserted from the inside.
      • We found that the coldhead RTD had failed during intial mounting efforts, and a new RTD (actually one that had been desoldered from the setup previously, removed from the workpiece on 2021.08.07 but found to have no issue) was soldered and attached to the cold head, under the spring clamp.
    • Each shield was reinstalled, including:
      • The newly cut slots were used to pass shields over the rigid copper bar.
      • Electrical cabling was threaded through the usual apertures.
      • The outer shield was positioned on the G10 spacers.
      • Rough alignment was completed based on clearance from the rigid copper bar, and line of sight to optical window.
    • Final touchups were implemented:
      • Aluminum foil covered unused outer shield apertures.
      • A small aluminum foil panel was placed underneath the rigid copper bar, to cover the slot in the inner shield.
      • Final clamping of the inner shield and the heater (with indium gasket) were completed.
      • Thermal strap was used to link the cold baseplate to the rigid copper bar, with a bolted joint at the copper bar and a dog clamp joint at the baseplate. Apiezon N grease was applied to all contact faces.

The rest of the installation effort is captured in the next log post QIL/2695, to partition the items relevant to the radiative cooling of the silicon mass.

The photos here (and others) are posted to the QIL Cryo Vacuum Chamber photo album.

Attachment 1: IMG_0352.JPG
IMG_0352.JPG
Attachment 2: IMG_0353.JPG
IMG_0353.JPG
Attachment 3: IMG_0371.JPG
IMG_0371.JPG
Attachment 4: IMG_0370.JPG
IMG_0370.JPG
Attachment 5: IMG_0343.JPG
IMG_0343.JPG
Attachment 6: IMG_0351.JPG
IMG_0351.JPG
Attachment 7: IMG_0358.JPG
IMG_0358.JPG
Attachment 8: IMG_0346.JPG
IMG_0346.JPG
Attachment 9: IMG_2734.jpeg
IMG_2734.jpeg
Attachment 10: IMG_2740.jpeg
IMG_2740.jpeg
  2693   Fri Nov 12 10:54:55 2021 ranaSummary2um PhotodiodesResults from JPL PD: A1-Test3

IT would also be good if you could plot the RMS noise around 10 Hz and 100 Hz as a function of the bias and temperature, so we can see what the trends are. And how about post the data and scripts to the elog so we can munge data later?

  2692   Fri Nov 12 08:21:22 2021 AidanSummary2um PhotodiodesResults from JPL PD: A1-Test3

[Aidan]

Here is the analyzed data from Test 3 of the A1 JPL PD.

Note about the QE measurement: 

  • HOM beam QE was performed during the warm up phase. Collimating lens was fixed and beam pointing was optimized at 100mA before each measurement. Likely that not all power was on PD but that distribution was constant throughout measurement. Therefore, good proxy for shape of QE response.
  • Manual QE was performed with 25mA current, optimized collimating lens position (and thus the beam size on the PD). The data corresponds to 8.0mm between the lens mount and fiber mount. The beam pointing was optimized before each measurement at 25mA.
  • QE projection scales the "HOM beam QE" result to the manual QE measurements to project out expected QE performance vs temperature

Dark current is the output from the Keithley scan - the vertical scale is correct in Amps [ignore question mark]

Dark noise spectra are included for different bias levels and at different temperatures. Stll need to add ADC noise floor for these plots.

Notes from Test 3

~/JPL_PD/data/A1_test3/README

6-Oct-2021: done with cryocooler off and temperature increasing

PREAMP GAIN = 1E3

SR560 gain = 500

LD temp set point = 20.2kOhm

 

 

 

Attachment 1: A1_test3_nominal_QE.pdf
A1_test3_nominal_QE.pdf
Attachment 2: A1_test3_darkcurrent.pdf
A1_test3_darkcurrent.pdf
Attachment 3: A1_test3_noise_spectra_rev.pdf
A1_test3_noise_spectra_rev.pdf A1_test3_noise_spectra_rev.pdf A1_test3_noise_spectra_rev.pdf A1_test3_noise_spectra_rev.pdf A1_test3_noise_spectra_rev.pdf
  2691   Wed Oct 27 10:18:33 2021 Aidan, StephenSummaryCryo vacuum chamberInspection of Megastat post 408K event and pumping timeline

[Aidan]

I've attached a timeline of our inspection this afternoon along with today's pumping timeline,

Here is a brief summary of observations from previous pumping timelines. Today's pump down is consistent with previously observed timelines.


 

2:17PM – Assessing the impact of the 408K event in the MegaStat

Coldhead reached 363K (90C)

Innershield reached 403K (130C)

 

2:24PM – Aquadag E service temperature (149C)

https://www.laddresearch.com/index.php/lanotattachments/download/file/id/40/store/1/

Maximum service temperature in air* : 300°F (149°C) 

*Service temperature under vacuum conditions is significantly higher. Contact Acheson for specifics. 

2:25PM – Bringing Megastat back up to air for initial inspection

2:37PM – chamber is at air

 

2:41PM – removing bolts

 

 

3:13PM – initial inspection looks normal. No elevated amount of black particulates found on surfaces – consistent with or less than the amount seen last time we opened.

Stephen detected faint smell different from last time (“campfire”?)

  • One RTD connector did delaminate Aquadag from the inner shield

3:14PM - Stephen reattaching the RTD wiring that had delaminated. I wiped up visible particulates with isoproponal soaked wipe.

 

3:21PM – putting lid back on

3:30PM – lid on. Screws in finger tight

 

3:35PM – screws tight – ready to pump

3:40PM – pumping station on

 

Time (minutes) Pressure (Torr) Notes
0 7.5E2 Pirani gauge initially
1 7.5E2  
2 7.5E2  
4 7.5E2  
5 7.5E2  
7 7.5E2  
9 7.5E2  
10 7.5E2  
11 7.5E2  
12 4.3E2 Gauge starts reading decrease
13 2.2E2  
14 9.8E1  
15 5.5E1 Turbo ON
16 2.9E1  
17 1.6E1 Turbo at 44%
18 9.4E0  
19 4.9E0  
20 1.8E0 Turbo at 58%
21 3E-2 Turbo at 70%
22 1.1E-3 ION gauge readings from here. Turbo at 91%
23 7.5E-4 Turbo at 100%
24 6.3E-4  
25 5.9E-4  
26 5.5E-4

 (Cryo-cooler normally turned on around this time)

Not in this instance though

27 5.0E-4  
28 4.59E-4  
29 4.26E-4  
30 4.05E-4  
     

 

Attachment 1: IMG_5407.jpg
IMG_5407.jpg
Attachment 2: IMG_5408.jpg
IMG_5408.jpg
Attachment 3: IMG_5409.jpg
IMG_5409.jpg
Attachment 4: IMG_5410.jpg
IMG_5410.jpg
Attachment 5: IMG_5411.jpg
IMG_5411.jpg
Attachment 6: IMG_5412.jpg
IMG_5412.jpg
Attachment 7: IMG_5414.jpg
IMG_5414.jpg
Attachment 8: IMG_5415.jpg
IMG_5415.jpg
  2690   Tue Oct 26 08:43:38 2021 AidanUpdateCryo vacuum chamberCTC100 temperature alarm and heater shutoff

Instructions on how to enable the alarm and heater shut off for the CTC100.

Status: This reports the status of the alarm. If LATCH is enabled, this must be manually set to OFF once it has been enabled.

Mode: Set to "Level"

Latch: Optional to set to "YES" if desired.

Output: Set to "Heater"

Max: Set to desired maximum temperature.

The attached photos show:

  • the menu where the settings are ALARM entered
  • the main display just before the alarm is enabled (at 300.350K with a 1s delay)
  • the main display just after the alarm is enabled - note that the Heater Output has been set to 0W.
Attachment 1: Screenshot_from_2021-10-26_08-42-57.png
Screenshot_from_2021-10-26_08-42-57.png
Attachment 2: IMG_5371.jpg
IMG_5371.jpg
Attachment 3: IMG_5396.jpg
IMG_5396.jpg
Attachment 4: IMG_5401.jpg
IMG_5401.jpg
  2689   Tue Oct 26 07:32:52 2021 AidanUpdateCryo vacuum chamberHeater left on - chamber got warm

We're at 300K as of 7AM this morning.

Quote:

The cryocooler was switched off last Thursday to do testing on the JPL_PD. I turned the heater back on during this testing and neglected to turn it off when I finished at the end of the day. As a result, the workpiece reached ~400K over the weekend.

We are now allowing it to slowly cool down.

The CTC100 has a feature to specify an upper limit on temperature and then shut off the heater if that temperature is exceeded. We should engage this going forward.

 

  2687   Mon Oct 25 16:13:47 2021 AidanUpdateCryo vacuum chamberHeater left on - chamber got warm

The cryocooler was switched off last Thursday to do testing on the JPL_PD. I turned the heater back on during this testing and neglected to turn it off when I finished at the end of the day. As a result, the workpiece reached ~400K over the weekend.

We are now allowing it to slowly cool down.

The CTC100 has a feature to specify an upper limit on temperature and then shut off the heater if that temperature is exceeded. We should engage this going forward.

Attachment 1: Screenshot_from_2021-10-25_16-19-39.png
Screenshot_from_2021-10-25_16-19-39.png
  2686   Mon Oct 25 16:10:51 2021 AidanUpdateCDSRestarted computers and front-end model following campus power outage

Rebooted the workstations and FB4.

 Restarted the model on the FB4:

  • sudo /opt/rtcds/caltech/c4/target/c4iop/scripts/startupC4rt
  • sudo /opt/rtcds/caltech/c4/target/c4tst/scripts/startupC4rt
  2685   Fri Oct 22 13:31:36 2021 ChrisLaser2um PhotodiodesImproved measurement of QE on photodiodes ~89% at 140K

Somewhere around the labs there should be a DataRay Beam'R2 scanning slit profiler, with an extended InGaAs detector that works out to 2.5 µm.

Quote:

agreed. You should put into Voyager chat and cryo/ET slack questins about 2 um beam profiling. We'll want it for anything 1.8-2.1 um.

Quote:

Definitely. I think the lack of beam profiling/imaging equipment is something we want to address too. We will waste a lot of time in Mariner if we can't profile our beams.

 

  2684   Fri Oct 22 12:04:13 2021 ranaLaser2um PhotodiodesImproved measurement of QE on photodiodes ~89% at 140K

agreed. You should put into Voyager chat and cryo/ET slack questins about 2 um beam profiling. We'll want it for anything 1.8-2.1 um.

Quote:

Definitely. I think the lack of beam profiling/imaging equipment is something we want to address too. We will waste a lot of time in Mariner if we can't profile our beams.

  2683   Fri Oct 22 09:20:13 2021 AidanLaser2um PhotodiodesImproved measurement of QE on photodiodes ~89% at 140K

Some of the data recorded during the current/micrometer scanning yesterday.

  • Distance between fiber/lens housings = Micrometer + 22.9mm
  • QE = Response / (1000Ohm*9.3E-4W/V) *(h*c/[e*L]) = Response * 0.667

Highlighted change to 25mA and also highest QE.

Time Temperature (K) Laser current (mA) Micrometer (mm) Peak Response (JPL/REF) QE
2:07PM 131.4 30 9.10 1.070 71.0%
2:14PM 133.5 30 9.10 1.090 73.0%
2:20PM 134.8 30 9.00 1.100 75.0%
2:25PM 136.4 30 8.80 1.134 75.6%
2:29PM 137.6 30 8.70 1.140 76.0%
2:31PM 138.1 30 8.60 1.120 74.7%
2:35PM 138.8 25 8.60 1.288 85.9%
2:42PM 140.5 25 8.80 1.246 83.1%
2:45PM   25 8.60 1.285 85.7%
2:249PM 142.2 25 8.40 1.310 87.4%
2:53PM 143 25 8.20 1.322 88.0%
3:01PM 144 25 8.00 1.328 88.6%
3:10PM   25 8.00 1.334 89.0%
3:14PM 147.5 25 7.80 1.314 87.6%
3:16PM 148.05 25 7.60 1.316 87.8%
3:19PM   25 8.00 1.328 88.6%
3:23PM 149.3 25 8.00 1.320 88.0%
3:30PM 151 25 8.00 1.315 87.7%
3:47PM 154 25 8.00 1.315 87.7%
4:02PM 157 25 8.00 1.300 86.7%
4:14PM 159.6 25 8.00 1.305 87.0%
  2682   Thu Oct 21 17:48:33 2021 AidanLaser2um PhotodiodesImproved measurement of QE on photodiodes ~89% at 140K

Definitely. I think the lack of beam profiling/imaging equipment is something we want to address too. We will waste a lot of time in Mariner if we can't profile our beams.

Quote:

its worth looking into how fiber optic mode cleaning actually works:

https://doi.org/10.1201/9780203739662

In order to get a lot of cleaning you have to have a clean beam to begin with. There's a way to pre-clean by putting the laser output into a pinhole before coupling int othe single-mode fiber. Also, use a ~40-50m fiber to make sure the mod-mistanatched beam actually goes into the cladding rather than recombine into the Gaussian beam.

 

  2681   Thu Oct 21 16:52:26 2021 ranaLaser2um PhotodiodesImproved measurement of QE on photodiodes ~89% at 140K

its worth looking into how fiber optic mode cleaning actually works:

https://doi.org/10.1201/9780203739662

In order to get a lot of cleaning you have to have a clean beam to begin with. There's a way to pre-clean by putting the laser output into a pinhole before coupling int othe single-mode fiber. Also, use a ~40-50m fiber to make sure the mod-mistanatched beam actually goes into the cladding rather than recombine into the Gaussian beam.

  2680   Thu Oct 21 15:32:32 2021 Aidan, RadhikaLaser2um PhotodiodesImproved measurement of QE on photodiodes ~89% at 140K

[Aidan, Radhika]

We turned off the heater and the cryocooler this morning to around 11:30AM when the temperature of the diode was around 123K and are letting it gradually warm up.

Through the next 30K, we experimented with different distances between the fiber output and the collimating lens. Bias voltage always set to 1000mV. Laser diode current was set manually on the controller (the input from the DAC was unplugged as this is a little noisy). 

  1. We increased the distance between collimator and the lens by 1.4mm (from stage reading of 9.41mm to 10.8mm) and there was a small increase in the response (PD output/REF PD output)
  2. For each new setting, we set the laser diode current to 100mA and run the maximize script on the piezo mirror, adjusting the alignment to maximize the output power.
  3. We then stepped up the laser diode current in steps of 10mA from 25mA to 95mA and one last measurement at 101mA. The PD response dropped by 30-40% through this range.

It looked like the optimum translation stage setting was 9.75mm - however, i discovered something very interesting ...

If you run the maximize power script at 100mA to the laser diode, then drop the laser current to 30mA and rerun the script, you find that there is a different optimum alignment. This means that the output beam shape/pointing is power dependent. In other words, the output of the fiber is not properly mode-cleaned by the 2m patch cord we have.

Switching to 25mA, I optimized the alignment and continued exploring the optimum translation stage position. Dropping the stage position to 8.0mm maximized the response (at 25mA). Note that the code maximizes an EPICS channel called C4:TST-PD_RESPONSE which is (JPL_PD - DarkV)/REFPD. The reference PD filter bank has an offset applied so OUT16 has a mean value of zero when the laser is off. DarkV for the JPL PD is the PD voltage when the laser is off and this was manually updated every 5-10 minutes or so. C4:TST_PD_RESPONSE is not stored but the JPL_PD and REFPD channels are stored in frames.

On 27-Sept, I measured the ratio of power incident on the JPL PD to the voltage output from the REF PD: dP/dV = 9.3E-4 W/V. The JPL PD DC photocurrent sees a transimpedance gain of 1000Ohm. Therefore, QE is calculated using the following formula:

QE = (RESPONSE)*(1E-3/9.3E-4)*  h*c/(e*lambda) = RESPONSE*0.667

Using this calculation and a peak response value of 1.334 at about 145K, the peak QE was estimated to be about 89%. An error analysis is needed on this. And we need to figure out how to get a better output beam shape from the optical fiber (use a really long fiber?)

Note that the translation stage reading of 8.0mm corresponds to a fiber holder to lens mount distance of 30.9mm

 

Attachment 1: IMG_5305.jpg
IMG_5305.jpg
  2679   Mon Oct 18 15:25:14 2021 AidanSummary2micronLasersStarting data taking and second test of JPL PD A1

Initial running of analysis code puts the max QE at ~62 + /- 1% around 130-150K. I want to explore this temperature regime manually and see if we're saturating the PD or not.

3:30PM - Chamber is still under vacuum. Cryocooler turned back on.

Quote:

Terminated the data taking at 8:35Am this morning. The termperature traces of the cryo chamber show a couple of discontinuities in the gradient. I don't know what the cause is,

Quote:
  • We're at 164K as of 8AM this morning.

 

 

  2678   Mon Oct 11 08:35:21 2021 AidanSummary2micronLasersStarting data taking and second test of JPL PD A1

Terminated the data taking at 8:35Am this morning. The termperature traces of the cryo chamber show a couple of discontinuities in the gradient. I don't know what the cause is,

Quote:
  • We're at 164K as of 8AM this morning.

 

Attachment 1: Screenshot_from_2021-10-11_08-36-06.png
Screenshot_from_2021-10-11_08-36-06.png
  2677   Thu Oct 7 08:07:03 2021 AidanSummary2micronLasersStarting data taking and second test of JPL PD A1
  • We're at 164K as of 8AM this morning.
  2676   Wed Oct 6 13:50:18 2021 AidanSummary2micronLasersStarting data taking and second test of JPL PD A1
  • Turned cryocooler off around 1317588441 (about 1:46PM)
  • Restarted measurement with output going to JPL_PD/data/A1_test3
  • Room is noticeably quieter without the cryocooler on.
Quote:
  • Output going to JPL_PD/data/A1_test2 and DAQ
  • Test commenced at 8:20AM and cryo cooler started shortly afterwards
  • Once trhough the loop takes about 20 minutes
  • Cryocooler on at 8:42AM

 

  2675   Sun Oct 3 08:22:49 2021 AidanSummary2micronLasersStarting data taking and second test of JPL PD A1
  • Output going to JPL_PD/data/A1_test2 and DAQ
  • Test commenced at 8:20AM and cryo cooler started shortly afterwards
  • Once trhough the loop takes about 20 minutes
  • Cryocooler on at 8:42AM
  2674   Sat Oct 2 23:57:50 2021 StephenDailyProgressCryo vacuum chamberChamber pumping down, carbon paint flakes cleaned up

Pump down sequence executed tonight; Aidan plans to automate data collection during cooldown and warmup both, and the script will be activated early in the coming week.

- Carbon paint flakes (mentioned by Aidan in QIL/2667) were either picked up or scrubbed by IPA wipe, except the biggest, which were nudged near the closest 1/4-20 hole and picked up with tweezers for removal. There are still some small flakes of paint as there was less benefit to cleaning flakes closer to the PD or lens, and I opted not to risk any bumps. Aidan was correct, some areas of inner shield ID wall have flaked, but it seems the main location of delaminated paint is on the wrinkly foil excess covering the rim of the shield, an accidental paint location.

- While adding the lids, I monitored the PD outputs using Aidan's strip tool kindly left running. I never noticed any clipping in the trends - should I have been more skeptical?

- While adding the lids, I forgot to monitor the RTD outputs to the CTC100 controller, and the outer shield ended up shorting and ceasing to read any temperature. Didn't notice until I had turned on the roughing pump! Had to reopen and fix the short. Good reminder.

- Turbo came on at 12:07 am on 03 October at a pressure of 30 Torr - the setting is actually a timer and not a gauge reading.

  2673   Fri Oct 1 17:41:25 2021 PacoMiscEquipment transferFiber alignment pen

We borrowed the fiber alignment pen from the QIL for quickly coupling the AOM first-order beam in the ECDL experiment in DOPO lab.

  2672   Tue Sep 28 08:23:11 2021 AidanUpdate2um PhotodiodesDistances between optics, collimating lens, focussing lens and photodiode

Precise distances required between:

  • fiber launcher and collimating lens
  • focussing lens and beam waist 

accounting for thickness of optic mounts, sunken fiber launcher plane, back focal length of lenses, dispersive variation in focal lengths of lenses  from nominal and distance between PD surface and base of PD mount. Also shown are the distances between the steering mirrors (PZT steering mirror, lower periscope mirror and upper periscope mirror).

Beam propagation through this system is shown in the attached PDF. The upper plot shows a paraxial beam propagation as the collimating lens is displaced from the nominal position. The purpose is to indicate the beam size (radius) all the way through the system. We would like this to be less than about 6mm radius (12mm diameter) on all of our 1 diameter optics. The second plot shows the waist size at the PD as the collimating lens is moved by +/- 2mm. The purpose is to allow us to tune the beam size on the PD without clipping the beam on intervening optics. 

Keeping the collimating lens Delta Z to a range of +/- 2mm is safe for beam propagation in terms of clipping on apertures or on the 1.5mm diameter PD.

Attachment 1: JPL_PD_collimating_lens_Optical_layout.pdf
JPL_PD_collimating_lens_Optical_layout.pdf
Attachment 2: JPL_PD_optical_propagation_and_beam_size.pdf
JPL_PD_optical_propagation_and_beam_size.pdf
  2671   Wed Sep 22 16:40:19 2021 AidanLaser2um PhotodiodesBeam size measurements of the 2um beam on the PD

I performed some occlusion measurements of the 2um laser going into the cryo chamber. For different values of dz on the collimating lens translation stage, I moved the power meter into the beam using it's translation stage by an amount dx.

One the beam was on the power meter (aperture = 5mm diameter) the power stayed constant for several MM before dropping again (indicating all the laser beam was on the power meter).

There was a big inrcease in incident power as dz was increased. This, and the constant power across the PD aperture, indicates that the beam is clipping or sees an aperture somewhere like the focussing lens (f=75mm) or further upstrean. I will review the expected beam size as a function of position, assuming the given NA fof the fiber.

 

Attachment 1: PD_measurement_layout.png
PD_measurement_layout.png
Attachment 2: PD_occlusion_measurements.txt
	Length from Focussing lens to POW METER = 80mm															
																
																
2.73mm			3.73mm			dz	4.73mm	5.73mm	6.73mm	7.73mm	8.73mm	9.73mm	10.73mm	11.73mm	12.73mm	13.73mm
																
dx (0.01mm)	PM(uW)		dx	PM(uW)		dx	PM(uW)	PM(uW)	PM(uW)	PM(uW)	PM(uW)	PM(uW)	PM(uW)	PM(uW)	PM(uW)	PM(uW)
60	191		40	3		200	241.9	272	305	348	395	454	523	605	697	798
50	189		50	7.25		175	241.7	270	307	346	394	452	519	595	685	777
35	181.7		65	23.6		150	234.5	262	296	333	377	432	492	563	645	727
20	169		75	85.5		130	218.8	243.6	274	308	356	395	445	507	579	657
... 7 more lines ...
  2670   Mon Sep 20 14:26:38 2021 ranaComputing2um PhotodiodesAutomation and analysis scripts for 2um data taking

you can put these in the GIT repo for the QIL Cryo tests that Radhika set up. Otherwise, they'll get lost. And we should probably change autorun to a .py script and document these in the README on the repo.

Quote:

The attached files are the scripts used to take data during the PD temperature cycling/testing and to retrieve and analyze data after the fact.

  • ~/JPL_PD/scripts/autorun2021.sh
    • ~/JPL_PD/scripts/piezo_mirror/maximize_output_power.py
  • ~/JPL_PD/data/A1_analysis/A1_analysis.py

 

  2669   Thu Sep 16 10:33:59 2021 AidanComputing2um PhotodiodesAutomation and analysis scripts for 2um data taking

The attached files are the scripts used to take data during the PD temperature cycling/testing and to retrieve and analyze data after the fact.

  • ~/JPL_PD/scripts/autorun2021.sh
    • ~/JPL_PD/scripts/piezo_mirror/maximize_output_power.py
  • ~/JPL_PD/data/A1_analysis/A1_analysis.py
Attachment 1: autorun2021.sh
#diode name
i=1001
diode=A1
caput C4:TST-FM15_OFFSET 0
sleep 1
while :; do
        #-----------------------------------------------------
        # dark current
        echo =======================
        echo ----- TOP OF LOOP -----
... 141 more lines ...
Attachment 2: maximize_output_power.py
# script to maximize the output power of the piezo
import serial
import time
import os, sys, subprocess
import numpy as np

def slowDownJog(ser):
    ser.write('1SU50\r\n')
    time.sleep(0.1)
... 195 more lines ...
Attachment 3: A1_analysis.py
# analysis od the A1 JPL PD diode
# Aidan Brooks - 10-Sept-2021

import cdsutils
import numpy as np
import matplotlib.pyplot as plt
import os, glob
import scipy.signal

... 172 more lines ...
  2668   Wed Sep 15 08:33:39 2021 AidanUpdate2um PhotodiodesVideo review of 2um testing setup post A1 testing

https://dcc.ligo.org/LIGO-G2102040

 

  2667   Wed Sep 15 08:22:32 2021 AidanDailyProgressCryo vacuum chamberCONTAMINATION: Black paint flecks throughout chamber

I was setting up for some characterization measurements of the JPL PD and I noticed that there are flecks of black paint all through the chamber. There were a couple of visible bare sections on the wall of the inner shield where paint had been removed.

Quote:

Monday I completed the vent that Aidan had started by turning off the cryocooler. During the afternoon I turned off the pumps, unbolted the chamber lid, and removed the radiation shield lids.

Next, Aidan was going to run some characterization measurements and determine whether to swap the diode or repeat with A1.

 

Attachment 1: IMG_4686.jpg
IMG_4686.jpg
Attachment 2: IMG_4687.jpg
IMG_4687.jpg
Attachment 3: IMG_4689.jpg
IMG_4689.jpg
Attachment 4: IMG_4690.jpg
IMG_4690.jpg
  2666   Tue Sep 14 15:58:51 2021 RadhikaDailyProgressCryo vacuum chamberCold plate cooling limited by copper braid

Koji asked me to calculate the thermal resistance between the cold head and cold plate from Megastat cooldown data, to compare to the theoretical thermal resistance of the copper braid. This way we can determine if cold plate cooling is limited by the braid itself or by the contact(s) between the braid and cold head / cold plate. 

After folding the copper braid in half, its cross-sectional area is 1.34e-4 m2 (source here) and I estimated its length to be 30 cm. I used a room-temperature value for the thermal conductivity of copper, for simplicity (~400 W/m*K). The "theoretical" thermal resistance of the copper braid should therefore be 5.57 K/W.

I used existing cooldown data from the cold head and cold plate to fit the thermal resistance between the two. I ignored effects from room temperature and simply modeled conductive cooling from the cold head to the cold plate. The result of the fit was a thermal resistance of 5.75 K/W, obtained from data. This value is pretty consistent with the calculation above, implying that the cold plate cooling is hitting the physical limits of the copper braid.

If the copper strap were instead a solid bar with the same nominal diameter (0.483"), the thermal resistance would drop to 3.15 K/W (a factor of 0.57 in cooldown time). 

  2665   Tue Sep 14 15:17:03 2021 StephenDailyProgressCryo vacuum chamberChamber up to air, lids removed

Monday I completed the vent that Aidan had started by turning off the cryocooler. During the afternoon I turned off the pumps, unbolted the chamber lid, and removed the radiation shield lids.

Next, Aidan was going to run some characterization measurements and determine whether to swap the diode or repeat with A1.

  2664   Fri Sep 10 20:13:12 2021 AidanDailyProgress2um PhotodiodesRunning automated testing suite on A1 PD

I terminated the data taking around 5PM when the photodiode was at about 4C (277K).

 

Quote:

9AM: At 232K this morning at 9AM. Turned on the heater to 1W around 9:07AM to speed up the return to room temperature. 

10AM - set heater to 7W

11:50AM - At GPS = 1315334850 - Set heater to 20W. Also noticed that SR560 output was railing so set gain to 500

Quote:

We're at 170K as of 9AM this morning. At the current rate, we should reach 273K tomorrow morning. 

 

 

  2663   Fri Sep 10 09:11:05 2021 AidanDailyProgress2um PhotodiodesRunning automated testing suite on A1 PD

9AM: At 232K this morning at 9AM. Turned on the heater to 1W around 9:07AM to speed up the return to room temperature. 

10AM - set heater to 7W

11:50AM - At GPS = 1315334850 - Set heater to 20W. Also noticed that SR560 output was railing so set gain to 500

Quote:

We're at 170K as of 9AM this morning. At the current rate, we should reach 273K tomorrow morning. 

 

  2662   Thu Sep 9 09:23:56 2021 AidanDailyProgress2um PhotodiodesRunning automated testing suite on A1 PD

We're at 170K as of 9AM this morning. At the current rate, we should reach 273K tomorrow morning. 

  2661   Wed Sep 8 14:35:22 2021 AidanDailyProgress2um PhotodiodesRunning automated testing suite on A1 PD

[Aidan]

I turned off the cryocooler and the A1 PD is slowly coming up to room temperature from ~53K (it's currently at 78K).

There is an automated script (autorun2021.sh) running to acquire data from the PD during this process (it is attached):

  1. Record dark current sweep with Keithley
  2. Record PD noise spectra (in DAQ) with different biases [-100, 0, 300, 600, 1000]mV. This involves steppnig the bias and waiting for 60s while fast AC data from the diode goes into the frames. Preamp setting = 1E3, SR560 setting is attached. Note that the SR560 photo shows an OVERLOAD light. This is only when the laser is on and the bias is high. When running the bias stepping up, the overload light only came on for a moment when the bias reached 1000mV - it stayed on for about 2-3s during this time. It's possible that the SR560 may saturate as the noise level increases with temperature. I might have to drop the amplification a touch.
  3. Record the QE for the PD for different bias settings. Bias is set to [-100, 300, 600, 800, 1000]mV. The laser current is turned on for 10s and off for 15s. The laser current output has a 0.15Hz pole on it so the current slowly ramps up and down. This allows us to record the reference PD and JPL PD levels for increasing intensities (as the PD is observed to saturate since the beam is close to a waist/focus on the PD).
  4. Get bright current sweeps with the Keithley for 20mA and 100mA going to the laser
  5. Run the maximize_output_power.py Python script that makes small changes to the piezo steering mirror to maximize the output of a CALC channel which is the ratio of the JPL PD to the REF PD (JPL_PD/REF_PD) - this removes first order power fluctuations from the alignment process.
  6. Dump all measurements into a directory identified by the GPS time and measurement number
    1. ~/JPL_PD/data/A1/<GPS_TIME>-<MEAS_NUM>/

Example output is attached.

Note - I was originally running the code with a manual realignment each time. I switched to maximize_output_power.py around the 2:30PM mark (~78K) and this yielded a 30% increase in photocurrent. So QE results below this are going to be low.

CALC channel (~/JPL_PD/Ioc/QIL/db)

record(calc, "C4:TST-PD_RESPONSE")
{
    field(SCAN, ".1 second")
    field(INPA, "C4:TST-FM29_OUT16")
    field(INPB, "C4:TST-FM30_OUT16")
    field(CALC, "ABS(B/A)<2?(B/A):2")
}

---

Notes from the lab:

Wednesday 8-Sep-2021

11:47AM – Script is ready to record data

  1. Record dark current sweep with Keithley
  2. Record PD noise spectra for different biases
  3. Measure QE for different biases
  4. Get photocurrent sweep vs bias with Keithley at 100mA and 20mA

 

12:15PM – start run with chiller on

  • 1315163811-1001 (first directory) – 53.6K

12:26PM- chiller off

  • 1315164407-1002 – 53.66K

12:36PM – third loop

  • 1315165029-1003 – 57.92K

12:48PM – 1315165720-1004: 60.98K

12:59PM – 1315166396-1005: 63.2K

1:10PM – 1315167064-1006: 65.2K

1:24PM – 1315167908-1007: 67.6K

1:35PM – 1315168569-1008 :69.4K

1:47PM – 1315169251-1009: 71.3K

  • Added a channel C4:TST-PD_RESPONSE which is PD_OUT/REF_PD. Cancels out the power fluctuations to first order to help when aligning.
  • This is running is an adHoc softIOC channel

1:57PM – 1315169873-1010: 72.9K

  • Might need to add heat into the base plate

2:07PM – 1315170489-1011 :74.45K

  • SR560 overloads a tiny bit on the bias = 1000mV setting for 3 or 4s after it reaches that setting.

2:26PM – running the max power script. Getting 30-40% more power!!

2:32PM – 1315171982-1001: 78.145K

NOW WITH AUTOALIGNMENT

2:34PM – 1315172074-1001: 78.3K

2:45PM – restarted code with longer pause (60s) between end of loop and maximize output power

3:18PM – restarted code with 120s pause between end of loop and Maximize Output

3:43PM – ADDED OFFSET OF 141 to FM29 (REF PD) which sets the zero power level to about zero. (in the middle of 1315176124-1003 measurement at 87.5K)

-----------

Channels:

Physical Variable Channel Name
Laser current setting C4:TST-FM12_OUT16
PD bias setting C4:TST-FM13_OFFSET
REF PD reading C4:TST-FM29_OUT16
JPL PD reading (DC) - preamp only C4:TST-FM30_OUT16
JPL PD reading (AC) - preamp + SR560 C4:TST-FM31_OUT
Keithley/Preamp switch C4:TST-FM15_OFFSET

 

Attachment 1: autorun2021.sh
#diode name
i=1001
diode=A1
caput C4:TST-FM15_OFFSET 0
sleep 1
while :; do
        #-----------------------------------------------------
        # dark current
        echo =======================
        echo ----- TOP OF LOOP -----
... 139 more lines ...
Attachment 2: maximize_output_power.py
# script to maximize the output power of the piezo
import serial
import time
import os, sys, subprocess
import numpy as np

def slowDownJog(ser):
    ser.write('1SU50\r\n')
    time.sleep(0.1)
... 195 more lines ...
Attachment 3: 1315172753_conditions.txt
C4:CTC-MS_WORKPIECE_TEMP_VAL 79.933
C4:CTC-MS_OUTERSHIELD_TEMP_VAL 241.116
C4:CTC-MS_INNERSHIELD_TEMP_VAL 86.579
C4:CTC-MS_COLDHEAD_TEMP_VAL 129.619
spectra_start 1315172767
PD_gain_DC FM30 C4:TST-FM30_GAIN 2
QE_v_bias_start 1315173068
Keithley_bright_start 1315173226
C4:CTC-MS_WORKPIECE_TEMP_VAL 81.07
C4:CTC-MS_OUTERSHIELD_TEMP_VAL 241.188
... 3 more lines ...
Attachment 4: IMG_4627.jpg
IMG_4627.jpg
Attachment 5: IMG_4628.jpg
IMG_4628.jpg
Attachment 6: Screen_Shot_2021-09-08_at_3.05.19_PM.png
Screen_Shot_2021-09-08_at_3.05.19_PM.png
Attachment 7: Screen_Shot_2021-09-08_at_3.05.51_PM.png
Screen_Shot_2021-09-08_at_3.05.51_PM.png
  2660   Tue Sep 7 11:52:40 2021 ranaComputingCymacs+/-18V power supply to AI/AA chassis was actually +18V, -14.5V

We should not have a bench power supply installed permanently. Can you install a Sorensen in that rack or use one of the nearby ones?

  2659   Tue Sep 7 09:14:22 2021 shrutiMiscEquipment transferTED200C borrowed

I borrowed (retrieved?) the TED200C temperature controller from the north table in QIL to use in the cryo lab.

  2658   Tue Sep 7 08:12:50 2021 AidanComputingCymacs+/-18V power supply to AI/AA chassis was actually +18V, -14.5V

[Aidan]

I was working in the QIL on Friday and I heard a clicking sound coming from the rack where the DAQ is installed. It turned out to be the DC power supply for the AI/AA chassis. One of the voltage was floating around from ~14.2V to ~14.8V and the unit was clicking as it did this. Since the AA/AI chassis expect +/-18V which is regulated down to +/-15V, this was, to use the scientific term, bad.

I set the low voltage channel back to 18V. We have noticed previous drifts DAC channels - it's possible this was the cause.

 

Attachment 1: IMG_4578.jpg
IMG_4578.jpg
Attachment 2: IMG_4577.jpg
IMG_4577.jpg
  2657   Mon Aug 30 08:04:52 2021 AidanDailyProgress2um PhotodiodesMonday morning observation of JPL A1 PD cool down

Attached are time series of the A1 photodiode temperature (and megastat temperatures) along with the PD response and reference PD measurement of delivered power. The laser is being modulated by a ramp wave (sawtooth) so the response shows the maximum and minimum. The difference between the maximum and minimum trend lines is an indication of the PD response (although the PD is saturating as well at higher intensities so further data analysis is required).

 

Attachment 1: Screen_Shot_2021-08-30_at_8.04.15_AM.png
Screen_Shot_2021-08-30_at_8.04.15_AM.png
  2656   Sun Aug 29 11:17:43 2021 AidanDailyProgress2um PhotodiodesStart of PD pumpdown

Sunday morning update:

11:17AM - Temperature of photodiode is ~86K

Also shown is the photodiode current time series. The laser is being ramped from minimum to maximum every 50s. The span of the signal increases as temperature is decreased. The response is almost nil at room temperature. Then it peaks around 140K and is clearly reduced by the time we get below 100K. (Caveat: the beam was initially aligned onto the PD at room temperature but there is currently no auto-alignment. The diode is quite large, 1.5mm sq, but we can;t rule out that the beam is moving off the PD as temperature decreases - so we can only put an lower limit on QE vs temperature from this data).

Attachment 1: Screen_Shot_2021-08-29_at_11.16.55_AM.png
Screen_Shot_2021-08-29_at_11.16.55_AM.png
Attachment 2: Screen_Shot_2021-08-29_at_11.22.03_AM.png
Screen_Shot_2021-08-29_at_11.22.03_AM.png
  2655   Sat Aug 28 07:01:21 2021 AidanDailyProgress2um PhotodiodesStart of PD pumpdown

Note on the script - it's running in a TMUX session on WS1: On for 15s, Off for 15s. A much better way to do this would've been to turn on an EXC in the laser current channel (I only remembered the AWG once I'd left). I'll pop in this morning to check on things and see if I can change the modulation.

The purpuose of the laser modulation is to continually monitor the PD dark current level as well as the change in the nominal QE as a function of temperature. I might change it to a sawtooth wave - if the PD is saturating from too much intensity, we'll see a decrease in the QE each time the laser power increases. FYI - we're getting about 0.8mW onto the PD and it should be close to the beam waist (numbers to follow on that); 

The DAC outputs are really wandering around a lot. We should diagnose what's going on there as this makes the laser current set point very noisy. 

10:00AM Saturday update: Temperature of the workpiece (photodiode) is around 192K. I terminated the script and started an excitation RAMP on the laser diode current. The response of the diode has increased dramatically and I think I see signs of the QE rolling off as the incident power increases.

Quote:

[Aidan, Stephen]

After Aidan validated that model inputs were creating physical parameter changes, we proceeded with some last few checks before closing up. Notes:

 - Aidan set up a helpful script turning laser power on and off, and strip tool to follow PD and monitor PD signals. A strip tool chart was used to confirm that there was no loss of functionality or alignment during pump down.

 - Checked Heater function and Workpiece RTD response - all good.

 - Confirmed alignment by steering at periscope output mirror and watching PD voltage.

At this point Aidan gave a green light for pumpdown prep, and to start pumpdown and start cooldown. Notes:

 - Stephen disconnected the PD connector accidentally while trying to add strain relief (the irony is palpable). Reattachment of connector didn't seem to affect any signal levels.

 - During installation of radiation shield lids, shields became misaligned and PD signal fell (presumably due to clipping). Recovered previous signal levels by realligning outer shield.

 - Double checked that everything seemed good to go, no issues!

 - Timeline of pumpdown and cooldown:

     17:32 - Pumpdown started.

     17:47 - Turbo spin up started (pump station delay parameter).

     17:55 - Pressure dropped below 1 mTorr, so I started cryocooler.

     18:05 - Healthy so far - Pressure had come down half a decade more, and Coldhead RTD was reading 235 K.

 

Attachment 1: 20210828_-_IMG_4445.jpg
20210828_-_IMG_4445.jpg
Attachment 2: 20210828_-_IMG_4443.jpg
20210828_-_IMG_4443.jpg
  2654   Fri Aug 27 17:57:30 2021 StephenDailyProgress2um PhotodiodesStart of PD pumpdown

[Aidan, Stephen]

After Aidan validated that model inputs were creating physical parameter changes, we proceeded with some last few checks before closing up. Notes:

 - Aidan set up a helpful script turning laser power on and off, and strip tool to follow PD and monitor PD signals. A strip tool chart was used to confirm that there was no loss of functionality or alignment during pump down.

 - Checked Heater function and Workpiece RTD response - all good.

 - Confirmed alignment by steering at periscope output mirror and watching PD voltage.

At this point Aidan gave a green light for pumpdown prep, and to start pumpdown and start cooldown. Notes:

 - Stephen disconnected the PD connector accidentally while trying to add strain relief (the irony is palpable). Reattachment of connector didn't seem to affect any signal levels.

 - During installation of radiation shield lids, shields became misaligned and PD signal fell (presumably due to clipping). Recovered previous signal levels by realligning outer shield.

 - Double checked that everything seemed good to go, no issues!

 - Timeline of pumpdown and cooldown:

     17:32 - Pumpdown started.

     17:47 - Turbo spin up started (pump station delay parameter).

     17:55 - Pressure dropped below 1 mTorr, so I started cryocooler.

     18:05 - Healthy so far - Pressure had come down half a decade more, and Coldhead RTD was reading 235 K.

  2653   Fri Aug 27 15:33:28 2021 AidanComputingCymacs 

Got the DAC working by reactivating entries in the C4TST_cdsMuxMatrix.

  • FM12-DAC12 = 1 (laser current) works
  • FM13-DAC13 = 1 (bias) works
  • FM15-DAC15 = 1 doesnn't work, No output on BNC at the AI chassis

No problems with channels 12-14. However, channel 15 doesn't output anything at the AI chassis.

Using channel 14 on the AI chassis with FM15 input into it.

  • FM15-DAC14

 

  2652   Thu Aug 26 16:10:16 2021 AidanLab InfrastructureCleanlinessTwo under table equipment racks purchased for North Table - one installed so far

I bought two racks to house all our electronics which is currently taking up space on the North Table. The legs were exactly the same length as the clearance between the floor and the bottom fo the table - but not after I got the hacksaw involved. Thankfully the legs have plugs on the top whic cover up the sharp bits that I cut.

Next step is to start transferring the equipment below the table.

Attachment 1: IMG_4407.jpg
IMG_4407.jpg
Attachment 2: IMG_4423.jpg
IMG_4423.jpg
Attachment 3: IMG_4421.jpg
IMG_4421.jpg
Attachment 4: IMG_4422.jpg
IMG_4422.jpg
  2651   Thu Aug 26 14:31:46 2021 Aidan, StephenLaser2um PhotodiodesIncident power calibration of 2um PD in new cryo chamber

We're getting close to running this (2um beam is focussed onto the PD and we have piezo mirror steering and beam size tuning). However, all DAC channels are currently non-responsive. I'm going to rebooting the front-end.

Incident power calibration was performed with a S148C power meter head placed directly in front of the PD. For varying current levels, I recorded the pick-off PD voltage and the power meter reading.

Current (mA) REF PD (V) Power Meter (mW)
101.04 1.01 0.803
101.03 1.012 0.805
100.41 1.01 0.805
93.11 0.935 0.735
85.03 0.846 0.664
79.97 0.792 0.618
75.01 0.74 0.573
69.97 0.684 0.524
65 0.629 0.478
55.01 0.52 0.384
44.96 0.405 0.284
35 0.291 0.1834
25 0.176 0.0844
20 0.119 0.0348
0.03 0.079 3.00E-04

POW [mW] = 0.862[mW/V]*REF_PD [V] -0.067[mW]

Quote:

[Aidan, Stephen]

Worked toward aligning and characterizing beam on PD. Will complete next session.

Log:

Some difficulty aligning to the 2um beam, which is sensed by a thermal card. Aidan intends to upgrade with a fiber coupled visible laser, which could then be swapped interchangably for alignment.

The 1" mirror at the top of the periscope doesn't make sense, given larger apertures in shields and viewport. We looked for a nearby 2" replacement but did not have luck. Ended up swapping back in the gold-coated 2" mirror, even though it is thin enough to be a pain to mount.

Instrumented connector pins to DB9 pins using the following translation (ref Aidan's drawing for connector / PD pinout, ref drawing from QIL/2639)

            DB9   - 1 6 2 7 3 8 

    Connector - 2 3 4 5 6 7

 

 

Attachment 1: IMG_4420.jpg
IMG_4420.jpg
Attachment 2: REF_PD_POW_CAL_cryo.pdf
REF_PD_POW_CAL_cryo.pdf
  2650   Tue Aug 24 15:01:25 2021 StephenLaser2um Photodiodes2um PD in new cryo chamber

[Aidan, Stephen]

Worked toward aligning and characterizing beam on PD. Will complete next session.

Log:

Some difficulty aligning to the 2um beam, which is sensed by a thermal card. Aidan intends to upgrade with a fiber coupled visible laser, which could then be swapped interchangably for alignment.

The 1" mirror at the top of the periscope doesn't make sense, given larger apertures in shields and viewport. We looked for a nearby 2" replacement but did not have luck. Ended up swapping back in the gold-coated 2" mirror, even though it is thin enough to be a pain to mount.

Instrumented connector pins to DB9 pins using the following translation (ref Aidan's drawing for connector / PD pinout, ref drawing from QIL/2639)

            DB9   - 1 6 2 7 3 8 

    Connector - 2 3 4 5 6 7

 

  2649   Fri Aug 20 14:05:45 2021 RadhikaDailyProgressCryo vacuum chamberCooldown model fitting for MS

The fit parameters are 1. the emissivity of the test mass and 2. the emissivity of the inner shield. Turning on absolute_sigma=True, the covariance matrix is:

[[4.40872002e-07 1.98911860e-07]
 [1.98911860e-07 1.50052351e-07]].

Interpreting this, the standard deviation \sigma of the parameters is:

e_testmass (painted with Aquadag): 0.00066398.

e_innershield (rough Al): 0.00038737.

As discussed in today's meeting, these values much lower than expected. I'll look more into how scipy.curve_fit calculates these values, and will use the fitting script discussed to better quantify the error. 

  2648   Fri Aug 20 13:44:58 2021 ranaDailyProgressCryo vacuum chamberCooldown model fitting for MS

I think the asymptotic temperature in the model is missing the data. i.e. the steady state temperature should match up, but the recorded data terminates too soon. Probably should figure out what the missing term is.

Can you post the covariance matrix of your fit so that we can see what the fractional errors are on the physical parameters? (i.e. construct the fit function so that physical parameters which are unknown are the fit parameters.)

  2647   Thu Aug 19 14:34:10 2021 RadhikaDailyProgressCryo vacuum chamberCooldown model fitting for MS

Building on [2643], I realized that the conductive cooling term proportional to (T_{coldplate} - T_{testmass}) was also negligible. I added back in physical parameters for the 2 radiative terms (one from the cold plate, one from 295K) and used the emissivities of the test mass and inner shield as fit parameters:

F_e = (\frac{1}{e_{tm}} + (\frac{1}{e_{is}} - 1) \frac{A_{testmass}}{A_{innershield}})^{-1}

P_{rad} = F_e A_{testmass} \sigma (T_{coldplate}^4 - T_{testmass}^4)

F_{e295} = (\frac{1}{e_{tm}} + (\frac{1}{e_{is}} - 1) \frac{A_{testmass}}{A_{aperture}})^{-1}

P_{rad295} = F_{e295} A_{testmass} \sigma (295^4 - T_{testmass}^4)

I used Koji’s input to model radiative heating from 295K. I approximated the radius of the aperture (from which room temperature could be exposed) to be 3cm, and assumed radiative heat is emitted from this circle. 

The result of this fit can be seen in Attachment 1: [e_testmass, e_innershield] = [0.59736291, 0.20177643]. This would imply that Aquadag has an emissivity of about 0.6, and that the emissivity of rough aluminum is much higher than expected at 0.2.

I then used these parameters to model the cooldown, given: 1. both the test mass and inner shield surfaces are painted in Aquadag; 2. only the inner shield surface is painted in Aquadag. These models are shown in Attachment 2. 

Painting the inner shield in addition to the test mass would yield marginal improvement, as expected. However, painting the inner shield while removing Aquadag from the test mass would, according to this model, weaken the coupling further compared to the reverse case. This makes sense, since in Fe the effect of the inner shield’s emissivity is scaled by the ratio \frac{A_{testmass}}{A_{innershield}}, which is quite small. Increasing the emissivity of the test mass therefore makes more of a difference in the coupling. 

Attachment 1: model_fit_tm_painted.pdf
model_fit_tm_painted.pdf
Attachment 2: models_painted.pdf
models_painted.pdf
ELOG V3.1.3-